Product unit are produced at the rate of 3 units per useful direct labour hour. Direct labour idle time is 10% of hour paid for. Sales of 500 units are planned and finished goods stock is budget to rise by 40 units. Which of the following is direct labour budget (hours)

Group of answer choices

162

200 hours

2000

1620
Answer #1

Answer : 200 hours

Calculated as

Sale of units = 500

Ending units = 40

Production required = 500+40 = 540

Hours required = 540/3 = 180 hours

So total hours required = 180 * 100/90 = 200 hours